### v2

# Block: Balancing Load in LLM Serving with Context, Knowledge and Predictive Scheduling

Abstract.
This paper presents Block, a distributed scheduling framework designed to optimize load balancing and auto-provisioning across instances in large language model serving frameworks by leveraging contextual information from incoming requests. Unlike popular model serving systems that rely on monolithic and heuristic task schedulers, Block operates as a fully distributed, stateless, and predictive scheduling system to achieve low overhead, reliability, and scalability. It leverages the deterministic and predictable characteristics of LLM inference, such as host configurations, response lengths, and hardware performance, to make scheduling decisions based on accurately predicted metrics. Evaluation on a 12-GPU cluster shows that Block significantly outperforms heuristic schedulers, boosting serving capacity by up to 16.7% and reducing P99 tail latency by up to 49.5%. These performance gains remain consistent across diverse models, workloads, and configurations. Code and data are open-sourced.
1. Introduction
The rise of Large Language Models (LLMs) like GPT-4 (OpenAI et al., 2024), Llama  (Touvron et al., 2023), and Gemini (Team et al., 2024) has revolutionized modern applications, such as chatbots (OpenAI et al., 2024), virtual assistants  (Guan et al., 2023), code generation  (Jiang et al., 2024), and creative writing  (Gómez-Rodríguez and Williams, 2023). This places immense pressure on LLM inference serving systems, which must meet stringent latency requirements for a seamless user experience and to maximize throughput to handle growing demand (Shen and Sen, 2025; Cheng et al., 2025; Kwon et al., 2023; Yu et al., 2022). To this end, in recent years, we have observed significant progress in the development of LLM serving systems. Several key techniques, such as Continuous Batching (Yu et al., 2022), Paged Attention (Kwon et al., 2023), Chunked Prefill (Agrawal et al., 2023), and FlashAttention (Dao, 2023), have been developed with significant improvements in both latency and throughput.
However, the current LLM inference process is often characterized as unpredictable (Sun et al., 2024). For example, the autoregressive nature of LLMs, which generates tokens sequentially based on preceding ones until a stop signal is reached, leads to variable response lengths and decoding steps (Schuurmans et al., 2024). Paged Attention (Kwon et al., 2023), which dynamically allocates memory resources and allows for request preemption, further contributes to the dynamic nature of runtime memory consumption. Furthermore, latency of each decoding step also exhibits high variance, due to dynamic batch size.
The above uncertainties challenge both scheduling LLM requests and scaling model instances, since widely used runtime metrics such as latency, throughput and memory, no longer accurately represent end-to-end execution loads in the case of LLM serving. Further, scheduled requests with unexpectedly long responses can further increase memory load on overloaded hosts and block subsequent requests from being launched (Sun et al., 2024). Most current multi-instance serving frameworks in production, such as (nvi, [n. d.]; Aminabadi et al., 2022), employ scheduling heuristics like round-robin for request dispatching which offer no guarantee on scheduling performance. Alternative solutions, such as Llumnix (Sun et al., 2024), schedule tasks while aligning with dynamic re-balancing through live migration, demonstrating improvements in overall cluster serving performance. However, such solutions require transferring requests’ KV cache (Kang et al., 2024) across instances with extra cost of network bandwidth and memory, which may not be preferable in a resource-constrained cluster with high serving pressure.
Recent studies indicate that the uncertainty of inference behavior can be mitigated with the help of purposely built assistant models. For example, the length of responses and duration of execution can be accurately predicted with a pretrained regression model (Zheng et al., 2023b), and via the inference simulation framework (Agrawal et al., 2024a). This presents an unexplored opportunity to improve load balancing in inference systems by exploiting such approaches. For instance, consider the request to “explain the theory of relativity”, which involves short prompts but generates lengthy responses. Predicting a request’s length allows the scheduler to proactively route it to less-loaded devices, which would improve cluster balance.
In this paper, we present Block, a novel decentralized scheduler for LLM inference clusters. The key novelty of Block relies in leveraging queries’ context and utilizing static properties of LLM serving backends, such as hardware capability and serving framework batching strategy, to predictively schedule inference requests to serving instances.
Block operates in the following way. A lightweight LLM-based regression model is first applied to estimate the length of response based on the request contexts. Then, a simulation framework is applied to predict key target metrics for each request. Finally, the Block task scheduler dispatches requests based on these simulation-based predictions. Following the above steps, Block can estimate real serving load and metrics for each request during scheduling, thereby improving overall performance by maximizing both throughput and resource utilization while minimizing both latency and preemptions. Additionally, our work identifies that such simulation-based prediction provides an efficient approach for auto-provisioning.
We evaluate Block against widely-used heuristic dispatchers on a 12 GPUs cluster with real-world datasets and LLMs. Results demonstrate Block’s improved performance: it increases serving capacity by up to 16.7% and boosts throughput by up to 4.4% against the Llumnix dispatcher. Critically, Block reduces average request latency by 19.9-45.8% and P99 tail latency by 12.6-49.5% compared to baselines. Block’s performance gains are more pronounced when looking at the Time-To-First-Token (TTFT) where average and P99 TTFT are reduced by 88.1-97.0% and 78.6-94.5% respectively. Further, Block improves resource utilization and achieves a 20.1% reduction in P99 latency when using prediction for auto-provisioning.
In summary, this paper makes the following contributions. First, we thoroughly motivate our approach to using predictive scheduling for LLM serving with the integration of response length estimation and inference simulation techniques in §2. Second, we design (§4) and implement(§5) Block, a distributed scheduling framework with predictive scheduling. Third, we conduct a comprehensive evaluation on a 12 GPUs cluster with real-world datasets/models to demonstrate the superior effectiveness of Block on load balancing and auto-provisioning, as presented in §6. Finally, Block paves the way for a new class of predictive LLM schedulers for online serving. Block’s code and dataset are open-sourced at https://github.com/AKafakA/Block.
2. Background
In this section, we provide an overview of LLM inference development and cover key techniques, such as Continuous Batching, Paged Attention, and Chunked Prefill, that are foundational to our design.
LLM Inference
Modern large language models (LLMs) are mainly built on transformer architectures (Vaswani et al., 2017) to process input sequences. LLM queries typically consist of strings of variable lengths known as prompts or contexts. Queries are first tokenized and converted into a sequence of embeddings, which are then processed by transformer blocks (Vaswani et al., 2017) by applying attention mechanisms in conjunction with multilayer perceptrons to project them into a specified space. The procedure of processing prompts is known as encoding.
Most current LLMs are designed to generate output sequences, a phase commonly referred to as decoding. During decoding, LLMs follow an autoregressive approach to generate one token at a time, and each is used along with encoded prompts’ tokens as input for future decoding. The decoding process continues until either the last generated token matches a predefined end-of-sequence token, or the generated length reaches the maximum length constraint. Since the generated and context tokens are continuously reused, the intermediate key and value tensors that are used to calculate attention scores can be cached without recomputation. This technique is known as KV cache (Kang et al., 2024).
Due to autoregressive generation, the decoding phase cannot be accelerated through parallelization and requires caching all tokens until completed. The only exception occurs at the first decoded token, as it only uses the prompt as input without depending on a prior stage. This phase, which involves computing the KV cache for the context sequence, is referred to as the prefill phase, in contrast to the subsequent sequential decoding steps known as the decode phase. To this end, prefill is considered as compute-bound while decoding is regarded as memory-bandwidth-bound (Kamath et al., 2024).
Continuous Batching and Paged Attention
Modern GPUs can efficiently process large matrix manipulations, so requests are typically grouped into batches to fully leverage GPU resources. Each decoding step is designed to generate one token for all requests within the batch, and KV cache for all requests consumes increasing GPU memory due to the new token produced. Static batching was first applied by early inference frameworks  (Wang et al., 2021; Fang et al., 2021), which groups and processes a fixed set of requests. However, such a simple batching strategy is inefficiency, because requests often have variable response lengths. Processing slots for shorter, completed requests could get locked and wasted until the longest request in the batch finishes. To address this, Continuous Batching (Yu et al., 2022) enables completed requests to exit while new selected successors join in, usually in FCFS order, with a limitation on the maximum batch size. Finally, the local scheduler in inference frameworks is responsible for distributing incoming requests into execution batches.
Figure 1. Paged Attention and Preemption
A key challenge in Continuous Batching is memory allocation. Since response length is unknown at runtime, memory must be pre-allocated for the maximum possible sequence length and leads to memory fragmentation and wastage, especially for requests generating shorter responses. This inefficiency on memory utilization necessitates a more dynamic memory management solution. To address this challenge, Paged Attention (Kwon et al., 2023) was proposed and open-sourced through the vLLM serving framework. In the case of Paged Attention, instead of reserving maximum memory or risking an OutOfMemory error, vLLM divides GPU memory into fixed-size memory blocks and maintains a page table to track the locations of noncontinuous physical memory blocks for KV cache associated with each request. As illustrated in Figure 1, preemption occurs when memory is insufficient for the next decoding step. The newest request in the batch is preempted, returned to the head of the waiting queue, and its memory blocks are released for the remaining requests. Once memory becomes available, typically from completed requests, the preempted request resumes, requiring its KV cache to be recomputed. This dynamic memory allocation significantly improves the LLM serving instance’s capacity and throughput, making vLLM a leading framework in research and industry applications.
Chunked Prefill
A new key challenge that has recently emerged for LLM serving frameworks is efficiently scheduling and allocating memory for requests that involve both prefill and decoding, as mentioned above. The original vLLM scheduler often addresses this by creating separate batches for prefill and decoding requests, implementing a prefill priority strategy. This means prefill-only batches are created and executed as soon as new prefill requests arrive, potentially delaying or interrupting ongoing decoding batches. This approach effectively improves overall throughput and reduces TTFT. However, as shown in Figure 2, prioritizing prefill can interrupt ongoing decoding batches, leading to noticeable decoding stall bubbles and degraded tail latency.
Figure 2. Original vLLM and Chunked Prefill
To mitigate this trade-off, Chunked Prefill (Agrawal et al., 2023) with stall-free local scheduler is proposed (Agrawal et al., 2024b). This technique divides the prompt processing (prefill phase) into smaller, equal-sized chunks that can be executed across multiple scheduling steps. Hybrid batches are then formed, combining decoding steps with these prefill chunks. These batches operate under a defined token processing budget, interleaving decoding and piggyback the prefill chunks until the budget is exhausted.
Results (Agrawal et al., 2024b) confirm that Chunked Prefill significantly improves tail latency with only minor throughput reduction. Consequently, prominent serving frameworks like vLLM and SGLang (Zheng et al., 2024) use it as their default option. Furthermore, new execution kernels, such as PODAttention (Kamath et al., 2024) focus to optimize the execution of these hybrid batches. In addition to Chunked Prefill, the Prefill-Decode (P-D) disaggregation method (Zhong et al., 2024; Patel et al., 2024; Qin et al., 2024) mitigates interference between the prefill and decode stages by utilizing separate instances for each. However, it necessitates the extra transferring of KV cache from prefill instances to the decode instances.
3. Related Work and Motivation
Current LLM inference with dynamic batching and memory allocation as discussed above, introduces new cloud-based challenges for task scheduling and resource management. Accordingly, this section first elaborates on these issues and then summarizes applicable techniques, leading to the proposed solution discussed in §4.
Unpredictability in LLM Serving Scheduling
Figure 3 shows a typical architecture for LLM inference serving. Firstly, a user interacts with an API, which serves as the entry point to the inference framework. A local scheduler within the framework then batches these incoming requests for parallel execution. The LLM inference process is composed of a series of layers, primarily involving attention mechanisms and linear projections. To optimize performance, these operations are often accelerated by highly efficient kernel implementations (Dao, 2023; Ye et al., 2025). Besides, for real-world applications served on cloud (Wang et al., 2023; OpenAI et al., 2024; Enis and Hopkins, 2024), deploying multiple framework instances is typically necessary to ensure high availability under heavy request volume. This multi-instance setup requires an external global scheduler to distribute incoming requests effectively across the available instances for load balancing. While implementing a global scheduler may seem straightforward, it presents significant challenges due to the unpredictable nature of LLM inferences.
Figure 3. LLM Inference Systems
LLM inference systems are typically characterized for their high unpredictability on both the final generated results and overall system performance over time, as noted in (Sun et al., 2024). The unpredictability is primarily attributed to: 1) variable memory demands, resulting from unknown decoding lengths, which can trigger accidental preemption and lead to unexpected performance degradation; and 2) resource competition which can further cause interference between requests with runtime performance dropping.
This inherent unpredictability in resource requirements and execution behavior makes dispatching requests across LLM serving instances more challenging than generally dynamic task scheduling (Savvas and Kechadi, 2004). For example, a standard scheduler like Kubernetes might filter compute nodes based on initially available resources and then selects the node with the lowest current utilization. However, such static filtering and simple utilization metrics are often inadequate for LLM inference workloads because memory demands of requests are dynamic, varying with the decoding length. Additionally, the actual performance when executing a request can still be impacted by interference from other colocated requests (Sun et al., 2024). Furthermore, unexpected preemption can also occur leading to serving request reset and successors blocked.
These factors make it difficult to accurately predict resource demands and duration of a request on a given instance. Consequently, many existing dispatching strategies for routing requests across LLM instances rely on relatively simple heuristic approaches such as round-robin employed by systems like DeepSpeed-MII (Aminabadi et al., 2022) and Triton Inference Server (nvi, [n. d.]). Other frameworks take different approaches. For instance, vLLM often delegates the responsibility of cross-instance routing to the user or a higher-level orchestrator. In contrast, frameworks like SGLang implement their own heuristic-based routers that are tightly coupled with internal optimizations and features, such as RadixAttention. ServerlessLLM  (Fu et al., 2024) processes LLM requests as serverless functions and schedules them to instances with the least estimated model-loading startup time. LLM-d  (git, [n. d.]) serves LLM models in Kubernetes and supports P-D disaggregation. It features separate different customized schedulers for prefill and decode instances.
Such frameworks with rule-based heuristic schedulers could be carefully tuned to perform well under specific settings, e.g., configurations, workloads, or models. However, these schedulers are susceptible to performance degradation in dynamic environments and lack the quantifiable metrics needed to clearly explain how the internal parameters/rules impact user-facing metrics.
Auto-provisioning in LLM serving
To ensure a stable application and a good user experience in serving clusters with variable workloads, effective resource management must go beyond scheduling to include auto-provision to handle peak loads and guarantees performance under established SLOs. Although auto-provision is highly desirable and comes with well-known challenges often faced in other areas with similar scheduling characteristics, yet it is often overlooked in LLM serving systems design. Asynchronous cold starts in serverless computing (Liu et al., 2023) exemplify the issue. While a new instance is being provisioned and initialized, incoming requests continue to be routed to existing instances, exacerbating tail latency. The effect is even more pronounced for LLM serving. In addition to new arrivals, ongoing inference requests also continue generating tokens and consuming additional memory on already overloaded instances. This prolongs latency and increases load, even as the newly provisioned instances remain underutilized. Finally, the resulting load imbalance wastes resources and degrades serving performance.
Llumnix (Sun et al., 2024) can mitigate such imbalances through dynamic re-balancing. Initially, it dispatches requests using a heuristic scheduler. Subsequently, Llumnix performs continuous dynamic load rebalancing across instances by migrating active requests along with their KV caches. While dynamic rebalancing can mitigate runtime load imbalances, it requires significant GPU memory and inter-GPU network bandwidth to transfer associated token caches, impacting model parallelism performance. This resource contention can even intensify in either: 1) when auto-provisioning is triggered due to cold start issues, to re-balance load between heavily utilized existing hosts and newly provisioned ones; or 2) when other network-heavy features are applied, such as P-D disaggregation and tensor/pipeline parallelism (Shoeybi et al., 2020; Huang et al., 2019).
Despite the significant challenges caused by unpredictability in LLM scheduling and provisioning, recent research offers promising mitigation techniques that can improve scheduler design, which we explore below.
Offline Performance Simulator
With the rapid development and expansion of new LLM models, devices, and use cases, setting up an appropriate LLM serving cluster has become a critical yet challenging task for developers, particularly due to the high trial-and-error costs associated with GPU pricing. To address this, Vidur, the first LLM cluster simulation framework, has been proposed (Agrawal et al., 2024a). It aims to reduce potential hardware costs when searching for the optimal cluster configuration based on a given model, trace, and Service Level Objective (SLO) requirements. The insight behind Vidur is that since the local scheduler and batching logic are deterministic, and response lengths are typically available in the replay trace used to evaluate cluster performance, it is feasible to simulate the entire replay process if the execution time for each batch can be inferred. Additionally, by profiling low-level operators like attention and linear projection for a specific GPU, Vidur trains linear models to interpolate execution times for various batches. It finally achieves less than 9% error for key metrics such as throughput and latencies across a diverse set of models and GPU types. Besides, SimAI (Wang et al., 2025) also provides modeling and simulation for the entire LLM training process with average 98% accuracy.
Response Length Prediction
While some methods predict inference performance assuming known response lengths, alternative approaches also predict the initially unknown lengths to enhance inference itself. Recognizing that the variable length of responses is a key challenge, these techniques leverage prediction to improve local scheduler efficiency. For instance, Sequence Scheduling (Zheng et al., 2023b) proposes instruction tuning a smaller LLM specifically to predict the response length of the main serving LLM. This prediction allows their local scheduler to be optimized by grouping requests with similar expected lengths into the same batch. Such grouping minimizes computational waste and improves single instances’ throughput by 85%.
Similarly, LightLLM (Gong et al., 2025) predicts request output lengths using historical distributions. Based on these predictions, it infers the peak memory requirement for running batches, enabling the local scheduler to proactively avoid preemption. In the context of global scheduling, length estimation has also been utilized in recent works, primarily as a filtering mechanism prior to scheduling tasks. For instance, TetriServe (Hu et al., 2024) deploys a lightweight length-estimation model to filter out decoding candidate instances without sufficient GPU memory to accommodate the estimated number of tokens and then applies the Power-of-Two method (Richa et al., 2000) to select decoding instances with fewer pending requests. DynamoLLM (Stojkovic et al., 2024) classifies requests into three distinct pools based on their estimated lengths, scheduling them separately with associated host pools. To summarize, our insight is that by combining simulation techniques with length prediction, the unpredictability can be largely eliminated, which in turn creates an opportunity to improve multi-instance LLM serving performance.
4. Block System Design
In this section we present Block, a predictive task scheduler to fully leverage query context and cluster knowledge for load balancing and auto-provisioning. Block exploits our key insights from §2, along with the feasibility of mitigating unpredictability through length estimation (Zheng et al., 2023b; Gong et al., 2025; Hu et al., 2024) and simulation on LLM inference such as (Agrawal et al., 2024a; Wang et al., 2025).
Figure 4. Block Architecture
Block, as shown in Figure 4, comprises four services: query length tagger, global scheduler, predictor, and inference framework backend, to handle request distribution and results collection. The query length tagger service is the entry point of the system and is designed to predict and label the anticipated response lengths to requests. LLM Requests are executed on the GPUs running under an inference framework such as (Kwon et al., 2023; Zheng et al., 2024; Gong et al., 2025). The global scheduler is tasked with making scheduling and auto-provisioning decisions and directing requests to appropriate model instances to balance the load. When a new query is received, the global scheduler first disseminates prediction requests to the predictor services in order to collect instance statuses/predictions for scheduling. Then, it dispatches requests to the selected model instances. Further, the predictor service performs as a sidecar to predict the target metrics. The Block framework is designed to be agnostic to models, hardware, inference frameworks, and scheduling strategies.
4.1. Model Instance
A model instance is collection of services deployed on a GPU host that is responsible for model execution and response generation. It consists of two primary components: the main services of the inference framework and a sidecar service called the Predictor. As discussed in §2, different inference frameworks have been widely explored and developed such as (Kwon et al., 2023; Zheng et al., 2024; Hu et al., 2024; Yu et al., 2022; Aminabadi et al., 2022). Similar to Vidur, Block is designed to be a framework-agnostic, supporting various frameworks with dynamic batching. Each inference framework used by Block can be integrated into it individually by supporting a new status API to export its internal status, such as request lists and GPU memory blocks, for metrics prediction.
The predictor service runs locally on each instance and aggregates runtime data exported by the status API and transforms it into a metrics map via the predict API, consumed by the global scheduler. The Predictor’s main role is to simulate and predict key performance metrics, such as end-to-end latency or TTFT, for incoming requests to be used by Block scheduling. In cases where a request’s actual decoded length exceeds its predicted length, the simulator dynamically adjusts the estimated length for prediction by using the monitored decode length plus an another 10 steps.
The Predictor’s simulator is adapted from Vidur (Agrawal et al., 2024a) in the following ways. We redesign the Vidur simulator for single-instance prediction and encapsulated it within the Predictor service. As illustrated in Figure 4, this simulation involves a two-stage process. First, a local scheduler simulator models the batching strategy for a given inference backend. Second, a linear model predicts the execution times for the batches generated in the first stage. Since the model is static and inputs are received on-demand from the status API, the Predictor service is stateless and replicable for single instances to reduce simulation-related overhead through parallelization. Furthermore, the Predictor service is designed with an extendable interface, which could be implemented by alternative simulation frameworks such as SimAI (Wang et al., 2025).
4.2. Block and Baseline Scheduler
The global scheduler service is designed to be fully distributed and stateless to ensure scalability in large clusters. As illustrated in Figure 4, rather than maintaining a cached, global table of instance statuses, the global scheduler calls the predict API to obtain real-time metrics and predictions for its scheduling decisions. Although this could introduce additional overhead into the end-to-end latency, we anticipate this impact to be minimal, as measured in §6. Besides, the simulation process for scheduling is computationally intensive and relies on data exported from the inference framework. So, we run Predictors locally on instances and leave global scheduler focus on dispatching only. This approach reduces the overhead associated with both computation and data migration between services. This distributed architecture offers several key advantages. It enhances scheduling efficiency through parallel operations and improves scalability and reliability when serving with large-scale clusters.
Furthermore, our design does not require scheduling techniques that depend on complex, instance-side functionalities, such as live migration for dynamic rebalancing (Sun et al., 2024). These methods would necessitate centralized orchestration by the scheduler, a task beyond simple dispatching. Therefore, a fully distributed and stateless scheduler is sufficient.
4.3. Query Length Tagger
The query tagger service employs a lightweight proxy model to estimate responses’ lengths based on the input prompt and the serving model. This proxy is designed to operate as an online service in parallel, incurring minimal overhead. The architecture is pluggable, allowing for alternative estimators such as the model-free, sampling-based approach from LightLLM (Gong et al., 2025). If an incoming request already specifies a response length, or if a heuristic scheduler that does not require metric predictions is used, the query is forwarded directly to a randomly selected global scheduler. The estimated length is used solely for metrics prediction and then scheduling and does not affect the inference outputs.
5. Implementation
Block is built upon Vidur’s repository, extended with new modules that define the online services described in §4. All services are implemented with FastAPI (tia, [n. d.]) to align with various inference systems’ frontends (Kwon et al., 2023; Zheng et al., 2024; Gong et al., 2025).
Simulation-based prediction
Block’s Predictor service can use all six local scheduler simulators in Vidur including Sarathi-Serve, vLLM, and LightLLM, as well as linear models for batch latency predictions for simulation-based prediction.
Our performance analysis of Vidur’s simulation identified significant inefficiencies attributable to object duplication and suboptimal list operations (such as using list.pop(0)). While these bottlenecks were acceptable for Vidur’s intended offline use, they could introduce critical scheduling overhead for Block real-time predictions. To mitigate this, we re-implemented the primary simulation functions in Vidur and integrated a caching mechanism into the predictor. This cache memoizes latency predictions for previously seen batch configurations (defined by batch size and token count), substantially reducing the computational cost of the simulation.
Framework Integration
Block is designed to be backend-agnostic, decoupled from specific backend framework integration. The current Block prototype works with the integrated vLLM 0.7.2 base version. Integrating a backend involves two main steps. First, the backend’s internal state must be exposed via the new status API as discussed in §4.1. Second, the framework’s local scheduler simulator needs to be implemented to simulate the backend batching strategy. Both steps require minimal effort; for instance, the new vLLM simulator is only 161 lines of code (LoC), and the vLLM API commit is 154 LoC. Besides the integration, to reduce API overhead which could decrease simulation accuracy, we enable vLLM’s multi-process frontend, which separates the inference engine and API frontend into distinct processes. However, parsing dynamic JSON messages between services still incurs constant scheduling overhead (as presented in §6.3), due to Python’s GIL constraints on thread-level parallelism. To further mitigate this overhead, the framework could be migrated to an alternative RPC protocol like gRPC (grp, [n. d.]), or it could leverage Python’s GIL-free features once they become available in FastAPI and vLLM.
Global Scheduler Implementation
Block’s global Scheduler is designed to be highly flexible. Its metrics and strategy applied for scheduling are both configurable, enabling easy implementation of additional scheduling strategies based on single or multiple metrics predictions. For the current evaluation, we implement a prototype scheduler that selects the instance with the lowest predicted latency. For comparison, we also implement the following baseline schedulers within the same Block framework.
  • Random: randomly picks one instance without any context as input;
  • Round-Robin: schedules in a round-robin fashion, which is widely used by multiple production-grade model serving systems as  (nvi, [n. d.]; ray, [n. d.]);
  • Min QPM (Queries Per Minute): the default scheduling policy in LiteLLM (lit, [n. d.]), a popular open-source library for routing LLM requests in cloud environments. This policy simply selects instances with minimal latest QPM.
  • INFaaS++: the optimized version of INFaaS (Romero et al., 2021) implemented by Llmunix (Sun et al., 2024). The scheduling policy is simply defined as $usedMemory/batchSize$.
  • Llumnix-: refers solely to the improved heuristic dispatcher component of Llumnix (Sun et al., 2024), excluding its continuous rebalancing feature (see §3). Building on INFaaS++, it introduces a correction item by summing the required memory to prefill all pending requests as $prefillMemory$, to better measure the memory load with request prefill context. Its load is defined as $(usedMemory+prefillMemory)/batchSize$.
Length Estimation Model
In addition to the service and simulation functionalities, we also release the training data and scripts for the length estimation model. Consistent with other works with auxiliary models for LLM inference (Gong et al., 2025; Stojkovic et al., 2024; Hu et al., 2024), our model is designed to be lightweight, ensuring its overhead does not significantly impact end-to-end performance. While we initially considered the 7B model from Sequence Scheduling (Zheng et al., 2023b), its training and serving costs were prohibitive for us. We instead fine-tune a RoBERTa-base (Liu et al., 2019) regression model with 125M parameter, which offers a more efficient path to comparable performance. Evaluation is presented in §6.2.
Future Work
Our current prototype supports key components similarly to other LLM scheduling frameworks. As part of our future work, we plan to add features to align with key trends in LLM serving. While P-D disaggregation has not yet been integrated, we consider its support should be feasible. Similar to LLM-d (git, [n. d.]), it would involve dedicated schedulers for prefill and decode phases to enable inter-phase cache transfer. Given cache transfer is still under active development on inference frameworks, we defer the full implementation and detailed exploration of P-D disaggregation to future work. Additionally, cache transfer enables the exploration of Block Scheduler combined with dynamic rebalancing. Nevertheless, we anticipate that when integrating the above features Block’s performance advantages will persist, as the fundamental scheduling challenges discussed in §3 remain pertinent in disaggregated settings. Furthermore, incorporating prefix caching (vll, [n. d.]) can enhance simulator accuracy for real-world applications involving multi-turn conversations and repetitive prompts (Zheng et al., 2024). Since it has a negligible impact on experimental setup with de-duplicated dataset, we leave it for further exploration.
Current Block implementation comprises about 4,000 LoC and is released at https://github.com/AKafakA/Block with data and testing scripts.
6. Evaluation
We conducted a comprehensive evaluation of the Block scheduler against other baseline schedulers as described in §4.2. In this section, we first present the accuracy of the length estimation model and the backend simulator in §6.2. We then discuss the integrated end-to-end experiment results in §6.3 and analyze the underlying memory management behavior in §6.4. Results in §6.5 demonstrate that predicted metrics from online simulation can enhance resource provisioning. Finally, Block’s performance across varying models, configurations, and datasets are presented in §6.6.
Figure 5. Latency Prediction Metrics
6.1. Experimental Setup
Testbed
We evaluate Block on the CloudLab platform (Duplyakin et al., 2019) using 12 d7525 nodes each equipped with two 16-core AMD 7302 CPUs running at 3.00 GHz, 128 GB of ECC memory, one NVIDIA A30 GPU with 24 GB of memory, and a dual-port 100 GB NIC. In order to exploit all available GPUs for LLM online inference we execute the length estimation model at a dedicated host with a single L40 GPU to process the dataset in an offline manner. We make this adjustment solely for the current prototype and evaluation. According to our design, the length estimate model can seamlessly execute as an online service using a model serving framework such as TorchServe (pyt, [n. d.]).
Data, Configuration and Model
We use ShareGPT (sha, 2023), a real-world dataset consisting of 52K conversations sampled from chatGPT. During our experiments, the prompts in the conversation from ShareGPT are sent to the global scheduler in order following the Poisson distribution under varying arrival rates. We refer this rate as external QPS below.
When serving the model on vLLM, FlashInfer (Ye et al., 2025) is used as the attention library. Also, greedy decoding (Jurafsky and Martin, 2025) is used with temperature set as 0. The maximum request batch size is set to 48, and chunk size for Chunked Prefill as 512. This combination shows best performances among all tested configuration sets. To reduce the overhead caused by simulation, each host runs 16 Predictors in parallel to simulate and predict latencies. We observe such replication reduces total scheduling overhead up to 50% as described in §6.3.
Since each testing node in our testbed is equipped with a single GPU and limited bandwidth between nodes, co-serving across GPUs using tensor and pipeline parallelism  (Narayanan et al., 2021) is infeasible. So, we perform our evaluation using a medium-sized model that fits within a single node’s capacity. We select the LLaMA2-7B (Touvron et al., 2023), a widely used open-sourced standard LLM, with 16-bit floating-point quantization for the evaluation. The total model weight occupies 12.5 GB of GPU memory, split to 1056 memory blocks with vLLM’s default block size for KV cache. Furthermore, Vidur (Agrawal et al., 2024a) indicates that serving larger models with parallelism can reduce CPU overhead while increasing simulation accuracy, which in turn could lead to greater performance gains for the Block.
6.2. Prediction Accuracy
Table 1. Query Length Prediction
Metric Roberta Regressor Prompt-based LLM
Avg Error 78.755 62
Avg Error Rate 24.4% Not reported
Acc-50 69.93% 59%
Acc-100 77.15% 81%
(a) ACC-X: percentage of data points with error less than X.
6.2.1. Length Prediction
To train and evaluate our length prediction model, we divide the ShareGPT dataset into 40k training and 10k evaluation samples. We use tested model to generate responses and recorded the actual lengths, which serves as labels for fine-tuning the RoBERTa-base model on a host equipped with a L40 GPU.
The RoBERTa-base regression model (125M parameters) achieves accuracy comparable to the 7B prompt-based model reported in Sequence Scheduling (Zheng et al., 2023b) (presented in Table 1(a)), resulting in average 24.4% errors compared with labels. The Sequence Scheduling study also derives its metrics from 10k conversation requests, sampled from another conversation dataset (Zheng et al., 2023a). The offline evaluation of our 10k requests using PyTorch completes in only 4.8 seconds, demonstrating that potential online serving overhead from length estimation can be negligible. We use all 10k requests, tagged with both their real and estimated lengths for the evaluation. While this model was trained on a limited dataset, its performance is sufficient for our purposes. Also, its accuracy can be further improved significantly in a production environment with extensive, real-world data.
Figure 6. Request Metrics Under Different QPS
6.2.2. Simulation-based Metric Prediction
First, we assess the accuracy of the runtime simulator in predicting metrics and its effectiveness for scheduling, as shown in Figure 5. In this experiment, Block process requests with real length under varying QPS from 20 to 36 using a random scheduler, recording both predicted and actual serving latencies. Online requests have a 1% probability to be sampled. When triggered, a sampled request initiates two actions: 1) reporting the accumulated prediction error for all preceding requests, and 2) broadcasting itself to all instances and marking the instance with minimal predicted latency as selected.
Graphs in the top row show that the prediction error rate is stabilized and is consistently lower when using Chunked Prefill compared to prioritized prefill. This shows the effectiveness of Chunked Prefill in mitigating the disruptive stall bubbles that affect latency prediction. Further, overall error rates between 10% to 15% confirm the simulator’s high accuracy. This 10-15% prediction error is not expected to entirely skew instance selection, as it includes a constant overhead that uniformly impacts all instances and therefore does not alter their relative latency rankings.
Scatter plots in the middle row show serving latencies for sampled requests across all instances, demonstrate the correlation between predicted and actual latencies. These findings are further supported by the graphs in the bottom row which present the rank distribution of the same selected instances as above. Results indicate a high probability (40% to 80% with increasing QPS) that the scheduler selects the best-performing instance.
Figure 7. Average/Variance of GPU memory Blocks and Total Number of Preemption Under different QPS
[…]
These time-related figures have been smoothed using a unified gaussian filter to reduce noise and enhance readability.
6.3. Request Latency Performance
Figure 6 assess Block’s effectiveness on task scheduling. It compares Block against different baselines, namely random, round robin, Min QPM, INFaas++ and Llumnix- as described in §5) at varying external QPS.
We present results against the following performance metrics: 1) end-to-end (e2e) request latency as measured from benchmark clients; 2) TTFT as the duration from request arrival at vLLM to first token generation; 3) scheduling overhead defined as the difference between the end-to-end latencies and time spent at vLLM side; 4) capacity as the Max QPS Under SLO; following Vidur (Agrawal et al., 2024a), capacity is defined as the maximum QPS meeting a predefined SLO. We set this SLO to TTFT P99 ¡ 3 seconds, as this tail latency is highly sensitive to load, and performance degrades rapidly once this threshold is exceeded, as presented in Figure 6; and 5) request throughput defined as the number of requests divided by total experiments time.
Further, we also explored Llumnix v0.1.0 to enable full comparison against its dispatcher and live migration-based rescheduling. Although runnable, we encountered significant performance degradation when migration was enabled and the API client was crashing. We attribute this to our single-GPU-per-host setup, which lacks the high-speed intra-node communication, forcing reliance on slower inter-node RPC for KV cache migration between instances. This highlights previously unconsidered hardware-dependent costs and barriers for Llumnix. We therefore do not include Llumnix in the current evaluation.
We use Block* to denote Block operating with predicted lengths from our POC model for latency prediction. As in real-world applications, the actual prompt length could be available by prompt cache (Bang, 2023), with duplicated prompts and estimated lengths are only required for fewer new prompts. To measure single-precision capacity, we conduct further granular search around integer QPS bracketing SLO.
As shown in Figure 6, Block and Block* consistently outperform baseline schedulers across nearly every metric. They achieve the lowest mean and tail TTFT/e2e latency, along with the highest throughput. The CDF plots of TTFT and e2e are shown in Appendix A. Block shows additional overhead latencies (approximately 80 ms within capacity) compared to baseline schedulers. This extra overhead is primarily utilized by simulation for metrics’ predictions, whereas in other schedulers, overheads are attributed to data transferring and parsing. Besides, since predictors run in parallel, the overheads are independent of the cluster’s scale. Instead, it depends on the maximum waiting queue size across the instances and can increase linearly once capacity is exceeded.
Block* slightly underperforms compared to Block due to the error of length estimation but with less overhead addition. This discrepancy is due to the greater uniformity of the estimated output length relative to the actual length, causing higher hit rate on cached batch latencies during simulation. The overhead is below 3% of the e2e latency and tends to decrease once capacity is reached, as e2e latency then spikes more rapidly. Furthermore, INFaaS++ outperforms three basic schedulers under low QPS, but exhibits significant performance degradation with QPS increasing, particularly in tail latencies. Llumnix- mitigates this issue and outperforms other baselines by applying prefill length of pending queries as correction items over INFaaS++ load calculation as detailed in §5. Taking QPS 32 as an example, Block/Block* reduce average/P99 TTFT by 88.07%/78.6% and 23.58%/10.84% respectively. For e2e latency, Block/Block* achieve reductions of 19.87%/12.56% on mean and 3.55%/0.82% at tails. These improvements ultimately lead to throughput gains of 4.44%/2.53%. Similar trends hold across QPS and get more pronounced at higher QPS.
6.4. GPU Memory Utilization
The probed free memory blocks are not only required for Llumnix- and INFaaS++ dispatchers but also aid in investigating memory management behavior. We modify vLLM to export the cumulative number of preemptions. Results are shown in Figure 7, which are smoothed by gaussian filter to enhance readability. Plots in the first row refer to the average number of free blocks across instances before each scheduling followed by its variance as a measure of balance in the second row. The third row reports the accumulated preemption numbers for the cluster across incoming queries.
Results show that Block effectively balances memory usage across the cluster and can explain why comparison schedulers exhibit degraded performance. They tend to maintain high variance in GPU resources across instances, leading to preemption even when cluster-wide memory is sufficient. Further, when QPS is low, schedulers tend to focus on speeding up tasks, resulting in low variance and more free GPU memory blocks, as requests complete faster and preemption costs are thus avoided. When QPS increases beyond cluster capacity, resource limitations emerge and schedulers shift their priority to maximize resource utilization. This aims to prevent performance degradation caused by inefficiencies but results in frequent preemptions as memory more limited.
6.5. Auto Provisioning
Figure 8. Auto Provisioning with Different Strategy Under QPS=24
(a) Dotted lines at left-bottom figure indicate the change of scheduling instances when triggering provisioning.
As discussed in §3, provisioning based on runtime metrics could suffer from cold start. To tackle this, predicted metrics can be used to provision instances in a preemptive manner. To this end, we implement two simple auto-provisioning strategies in Block. The preempt strategy adds an instance when predicted latency reaches 70 seconds, while relief strategy means provisioning only when actual latency hits the same threshold. We conduct experiments starting with six available instances under a QPS of 24. We also establish a baseline with a sufficient static cluster of 10 instances.
Results in Figure 8(a) support our hypothesis that a relief strategy causes over-provisioning: as the newly added hosts are unable to relief queued requests, which could trigger further provisioning, all backup instances are quickly exhausted, and this causes memory imbalances and GPU resource wastage. The preemptive strategy activates earlier and lead to smoother changes in cluster size and ultimately provisioning only the necessary instances. This results in lower variance and higher utilization for memory overtime and reducing P99 latency by 20.1% (89.2/71.2) and requests over 70 seconds by 81% (627/129) with less instances.
6.6. Generality Study
Table 2. Scheduler Capacities with Setting Variables
Scheduler bs=24 cs=2048 qwen burstgpt
Block 27.9 31.5 68.3 59.0
Block* 27.2 30.8 67.9 /
Llumnix- 23.9 29.8 62 55.1
Gain 16.7%/13.8% 5.7%/4.3% 10.2%/9.5% 7.1%
(a) bs/cs means batch size and chunk size
Any dynamic changes on LLM inferences setting, as backend configuration, model, and data, are usually ignored by heuristic schedulers’ rules and could cause potential performance drift, as discussed in §3. In contrast, Block automatically integrates any changes of the serving cluster into simulations to fill the gap. We conduct a generality study comparing the capabilities of Block to Llumnix-, with different setting variables, as detailed in Table 2(a). Plots of other metrics and CDF of latencies are presented in Appendix B.
We experiment with sub-optimal configurations by varying the batch size or chunk size. Both changes lead to greater performance degradation on Llumnix- and enhance the advantage of Block/Block*, from the original 4.2%/1.3% gains in §6.3 to 16.7%/13.8% and 5.7%/4.3%. Then, we replace the model with Qwen2-7B (Yang et al., 2024) or data with BurstGPT (Wang et al., 2024), both generate shorter responses and lead higher capacity and demonstrate Block’s greater advantages. When testing with Qwen2-7B, capacity improves to 10.2% and 9.5% for Block/ Block*. Block* cannot run with BurstGPT dataset, since it only provides length traces without actual prompts to estimate the output length. Block is tested by generating prompts based on traces and shows 7.1% gain.
7. Conclusion
Block is a novel distributed scheduler that incorporates length estimation and simulation techniques to scheduling and auto-provisioning in LLM serving cluster. By one-shot, predictive scheduling, Block significantly boosts cluster capacity and reduces request latencies. Its predictive nature also allows for proactive auto-provisioning. Our work highlights the potential of predictive scheduling, paving the way for more efficient, responsive, and scalable LLM serving.
Appendix A Latency CDF for selected QPS
Figure 9. CDF for Latency Metrics
We present the CDF distribution for TTFT and end-to-end latency for selected QPS as 20, 24 and so on in Figure 9, which clearly confirm Block/Block* advantage on reduction tail latencies with high QPS.
Appendix B Supplementary Experiments with Generality Study
We presented all aggregated metrics and the CDF of selected QPS from Figure 10 to Figure 17 as below. For the Qwen and BurstGPT tests, we first applied binary search to roughly identify the wide range of QPS around capacity, as 48 to 64 for BurstGPT and 55 to 70 for Qwen experiments. We then conducted a granularity search from single integer to single float precision. As shown, the advantages of Block and Block* over Llumnix are consistent, leading to improvements in capacity, as summarized in § 6.6.
Figure 10. Metrics Under Different QPS for batch size = 24
Figure 11. CDF for batch size = 24
Figure 12. Metrics Under Different QPS for chunk size = 2048
Figure 13. CDF for chunk size = 2048
Figure 14. Metrics Under Different QPS for Dataset as BurstGPT
Figure 15. CDF for Dataset as BurstGPT
Figure 16. Metrics Under Different QPS for Model as Qwen2-7B
Figure 17. CDF for Model as Qwen2-7B